Bigger mammals tend to carry their young longer before birth. The length of horse pregnancies from conception to birth varies according to a roughly Normal distribution with mean 336 days and standard deviation 3 days.

� = � =

X- is ____________________________________ What percent of horse pregnancies are longer than 339 days?

With the questions marks one looks like an U the other question mark looks like an n
Answer #1

  Solution :

Given that ,

mean = \mu = 336

standard deviation = \sigma = 3

P(x >339) = 1 - P(x <339)

= 1 - P[(x - \mu ) / \sigma < (339-336) / 3]

= 1 - P(z <1 )

Using z table,

= 1 -0.8413

=0.1587

=15.87%

X- is _____________________15.87%_______________ percent of horse pregnancies are longer than 339 days